# DiffoRA: Enabling Parameter-Efficient LLM Fine-Tuning via Differential Low-Rank Matrix Adaptation

### v1

@@ 3.2 Main theorems
For an over-parameterized neural network with the loss on the testing set as $\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})$. Let $\bm{y}=(y_{1},...,y_{N})^{T}$, and $\eta=\kappa C_{1}\sqrt{\bm{y}^{T}(\bm{H}_{\Gamma,w_{0}}^{\infty})^{-1}\bm{y}}/(m\sqrt{N})$ for some small enough absolute constant $\kappa$, where $\eta$ denotes the step of SGD. Under the assumption of Theorem 2, for any $\delta\in(0,e^{-1}]$, there exists $m^{\ast}(\delta,N,\lambda_{0}^{\Gamma})$, such that if $m\geq m^{*}$, then with probability at least $1-\delta$, we have
@@
$\mathbb{E}[L(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})]\leq$ $\mathcal{O}(C^{\prime}\sqrt{\frac{\bm{y}^{T}\bm{y}}{\lambda_{0}^{\Gamma}N}})+\mathcal{O}(\sqrt{\frac{\log(1/\delta)}{N}})$ (9)
where $\lambda_{0}^{\Gamma}>\lambda_{0}$, $C,C^{\prime}$, and $\delta$ are constants.
Proof sketch. The proof of the above theorem derives from Corollary 3.10 of [1] and Section D.2 of [39]. We present the detailed proof in Sec. A.2.
Similarly to Theorem 2, the above theorem indicates that the adoption of $\bm{\Gamma}$ can enhance the generalization capability of the model, as shown in the last inequality.
Overall, in this section, we have proven that a selective matrix $\bm{\Gamma}$ will result in a higher convergence rate and better generalization capability of the network, thus leading to enhanced model performance. The main theorems in Theorem 2 and Theorem 3 theoretically grounded the effectiveness of the selective matrix. The question remains is how to construct $\bm{\Gamma}$ for the network efficiently. We will provide solutions and detailed algorithms in the next section.
4 Design of DiffoRA
We now present the concrete method to construct the module-wise selective matrix $\mathbf{\Gamma}$. The proposed approach is called DiffoRA, which is built atop the DAM (i.e., Differentiable Adaptation Matrix). To capture the information of the incremental updates modeled by the low-rank decomposition matrices, DiffoRA views the elements $\gamma$ in the module-wise DAM $\bm{\Gamma}$ as trainable parameters. More concretely, as shown in Fig. 1, DiffoRA approximates the NP-hard problem by invoking the following two stages: (i) Relaxation and optimization, which aims to map $\bm{\Gamma}$ to a continuous space (i.e., $\bar{\bm{\Gamma}}$) so that it is differentiable and can be further optimized; and (ii) Discretization and fine-tuning, which binarizes $\bar{\bm{\Gamma}}$ to determine the most essential module for fine-tuning. We will first elaborate on each stage respectively, and then introduce the weight-sharing optimization.
4.1 Continuous relaxation

### v2

@@ A.2 Proof of Theorem 2
Theorem 2. For an over-parameterized neural network with the loss on the testing set as $\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})$. Let $\bm{y}=(y_{1},...,y_{N})^{T}$, and $\eta=\kappa C_{1}\sqrt{\bm{y}^{T}(\bm{H}_{\bm{\Gamma},\bm{w}_{0}}^{\infty})^{-1}\bm{y}}/(m\sqrt{N})$ for some small enough absolute constant $\kappa$, where $\eta$ denotes the step of SGD. Under the assumption of Theorem 1, for any $\delta\in(0,e^{-1}]$, there exists $m^{\ast}(\delta,N,\lambda_{0}^{\bm{\Gamma}})$, such that if $m\geq m^{*}$, then with probability at least $1-\delta$, we have
@@
$\mathbb{E}[\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})]\leq$ $\mathcal{O}(C^{\prime}\sqrt{\frac{\bm{y}^{T}\bm{y}}{\lambda_{0}^{\bm{\Gamma}}N}})+\mathcal{O}(\sqrt{\frac{\log(1/\delta)}{N}})$ (21)
where $\lambda_{0}^{\bm{\Gamma}}\geq\lambda_{0}$, $C,C^{\prime}$, and $\delta$ are constants.
Proof.
According to the courant minimax principle [45], D.2 in [36], we get
$\bm{y}^{T}(\mathbf{H}_{\bm{\Gamma},\bm{w}_{0}}^{\infty})^{-1}\bm{y}\leq\frac{\bm{y}^{T}\bm{y}}{\lambda_{\min}(\mathbf{H}_{\bm{\Gamma},\bm{w}_{0}}^{\infty})}.$
Thus, we have
$\mathbb{E}[\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})]$
$\leq O\left(c\cdot\sqrt{\frac{\bm{y}^{T}(\mathbf{H}_{\bm{\Gamma},\bm{w}_{0}}^{\infty})^{-1}\bm{y}}{N}}\right)+O\left(\sqrt{\frac{\log(1/\delta)}{N}}\right)$
$\leq O\left(c\cdot\sqrt{\frac{\bm{y}^{T}\bm{y}}{N\lambda^{\bm{\Gamma}}_{0}}}\right)+O\left(\sqrt{\frac{\log(1/\delta)}{N}}\right)$
Method #Params MNLI SST-2 CoLA QQP QNLI RTE MRPC STS-B All
Acc Acc Mcc Acc Acc Acc Acc Corr Avg.
Full FT 124.65M 87.68 94.73 60.26 90.75 92.58 78.63 88.33 90.31 85.41
BitFit 0.1M 85.50 94.38 61.16 88.08 90.99 79.57 89.07 90.55 84.91
HAdapter 1.20M 86.53 93.73 62.62 90.83 92.82 80.43 89.90 90.16 85.88
PAdapter 1.19M 86.75 93.83 63.87 90.53 92.61 80.51 89.51 90.65 86.03
LoRA 1.33M 87.11 93.71 63.54 90.44 92.76 80.65 89.90 90.91 86.13
AdaLoRA 1.27M 87.89 95.11 63.23 90.48 92.84 81.23 89.02 91.22 86.38
FLoRA 1.33M 87.43 94.27 63.31 90.38 92.75 81.59 90.44 90.82 86.37
DiffoRA 1.32M 87.73 95.16 64.95 91.04 92.84 81.98 89.44 91.35 86.81
Appendix B Baseline details
We use the following methods as baselines. (i) Full FT uses all parameters for fine-tuning; (ii) BitFit [20] is a sparse-fine-tuning method for pre-trained models that updates only a small subset of the bias terms; (iii) Houlsby adapter [46] adds a few trainable modules inserted between layers of a pre-trained model, allowing for task-specific tuning without altering the entire model; (iv) Pfeiffer adapter [47] combines multiple task-specific adapters by linearly blending their outputs; (v) LoRA [9] reduces the number of trainable parameters by applying low-rank matrix decomposition to weight updates in pre-trained models; (vi) AdaLoRA [11] adapts LoRA by dynamically adjusting the rank of low-rank updates during training, optimizing parameter efficiency while maintaining model performance across various tasks. (vii) LoRA+ [27] employs different learning rates to update the low-rank matrices.
Appendix C Dataset details
Appendix D Training details
DiffoRA is fully implemented in Python. We evaluate the method on a Desktop Core i7-12700F CPU and Tesla V100 GPU. We summarize the training settings of GLUE and SQuAD in Table 6 and Table 7, respectively. In order to compare with the baseline method at the same parameter level, we use different $r$ for different datasets.
Appendix E Algorithm pseudocode
We present the training algorithm in Algorithm 1.